# LocoMan: Advancing Versatile Quadrupedal Dexterity with Lightweight Loco-Manipulators

### v1

@@ Abstract
Quadrupedal robots have emerged as versatile agents capable of locomoting and manipulating in complex environments. Traditional designs typically rely on the robot’s inherent body parts or incorporate top-mounted arms for manipulation tasks. However, these configurations may limit the robot’s operational dexterity, efficiency and adaptability, particularly in cluttered or constrained spaces. In this work, we present LocoMan, a dexterous quadrupedal robot with a novel morphology to perform versatile manipulation in diverse constrained environments. By equipping a Unitree Go1 robot with two low-cost and lightweight modular 3-DoF loco-manipulators on its front calves, LocoMan leverages the combined mobility and functionality of the legs and grippers for complex manipulation tasks that require precise 6D positioning of the end effector in a wide workspace. To harness the loco-manipulation capabilities of LocoMan, we introduce a unified control framework that extends the whole-body controller (WBC) to integrate the dynamics of loco-manipulators. Through experiments, we validate that the proposed whole-body controller can accurately and stably follow desired 6D trajectories of the end effector and torso, which, when combined with the large workspace from our design, facilitates a diverse set of challenging dexterous loco-manipulation tasks in confined spaces, such as opening doors, plugging into sockets, picking objects in narrow and low-lying spaces, and bimanual manipulation.
@@ I Introduction
Recent advances in the capability of quadrupedal robots have enabled them to traverse through highly complex terrains [1, 2, 3, 4, 5, 6], perform acrobatic skills [7, 8, 9, 10, 11], and interact with humans [12]. While most of these works focus on improving the mobility of quadrupedal robots, the manipulation capability of quadrupedal robots still remains limited and often requires special designs to achieve tasks such as button pressing [13] and ball kicking [14, 15]. To be effectively deployed in daily life, quadrupedal robots are required to possess versatile manipulation skills with enhanced dexterity, in addition to their inherent locomotion capability.
Integrating manipulation into quadrupedal robots proves to be a difficult task. As illustrated in Table. I, without external hardware modifications, quadrupedal robots have to utilize their legs, head or torso to move or transport objects [13, 15, 16, 17], which limits their ability to control the 6D poses of the object and perform high-precision tasks. A popular solution to address this challenge is to use a top-mounted robot arm [18]. However, this solution often comes at the cost of increased payload and decreased agility, and prevents the robot from reaching narrow spaces or performing more dexterous tasks such as bimanual manipulation. While foot-mounted grippers offer a lightweight alternative to full robot arms, existing works adopt 1-DoF grippers with limited functionality [19]. Therefore, it can be challenging to achieve versatile dexterous loco-manipulation using existing manipulator solutions for legged robots.
In this work, we present LocoMan, a comprehensive solution for versatile and skilled manipulation with legged robots. To reduce the payload and increase dexterity, LocoMan mounts two custom-designed, lightweight 3-DoF manipulators to the front calves of the robot, and uses the existing leg joints in addition to the manipulators for manipulation, as illustrated in Fig. 1. We refer to our designed manipulator as “loco-manipulator”, which draws conceptual inspiration from the front limbs of animals like great apes and bears, which are adept at both locomotion and skilled manipulation.
@@
    We demonstrate that LocoMan can perform versatile real-life manipulation tasks, including cabinet opening, charger inserting, drink pouring, grasping in narrow space, and basket hoisting.
@@ II-A Quadrupedal Loco-Manipulation
Without Additional Manipulators Researchers have had a long history exploring the manipulation capability of quadrupedal robots. Without an additional manipulator, researchers have designed the robot to use its head [17, 22, 23, 24], torso [16], or foot [13, 14, 15, 25, 26, 27, 28] to interact with objects in the environment, where [22, 23, 24] rely on multiple robots to push larger size objects. However, without an additional mechanism like a gripper, these approaches are limited to relatively simple tasks that do not require fetching objects, such as kicking balls [14], pressing buttons [13], pushing boxes [24], opening doors [27], and probing [26].
With Additional Manipulators Many previous works mount a dedicated arm on top of the quadrupedal robot to perform more complicated manipulation tasks such as pick-and-place [29, 30], pulling doors [31, 21, 32], tablecloth spreading [33], turning wheels [34], wiping a whiteboard [18], and collaboratively carrying large objects [35, 36]. On the other hand, equipping a robot with a dedicated arm typically adds to the mechanical complexity and power requirements, leading to an increase in both weight and expense. Moreover, such an addition makes it difficult to operate in a narrow space and reduce quadrupedal robot’s agility. Some other works alleviate these problems by installing more lightweight manipulators or grippers on the mouth [37] or foot [19, 20] of the quadrupedal robot. Tsvetkov et al. [19] propose a novel design of a small-scale quadrupedal robot with manipulators built into the legs, each requires three additional actuators, where two are used for grasping. It facilitates single-arm and two-arm manipulation but only supports specifying target positions in joint space. Arm et al. [20] propose learning-based controller that can track quadrupedal foot position targets through whole-body motions, and a 2-DoF gripper is attached to the foot to perform tasks like collecting rock samples. However, how to efficiently reach a specified 6D pose of the end effector on the leg to perform more complex manipulation tasks is still underexplored. While some hexapod robots [38, 39, 40, 41, 42] use legs as manipulators where the integrated grippers can fetch objects with high degrees of freedom, prehensile manipulation with dexterity could be challenging for quadrupedal robots.
@@ II-B Controllers for Leg-Manipulation
To accomplish a wide range of manipulation tasks, the controller of the loco-manipulation robot need to support a variety of operation modes, including reaching a narrow space with a single arm, walking while carrying an object, and bi-manual manipulation with upright pose. While prior works have achieved similar tasks such as grasping [21, 31, 18], object pushing and kicking [17, 14, 15], and object carrying [41] using optimization-based [21, 31, 29, 17, 41] or learning-based [16, 14, 15, 18, 13] methods, most of these controllers are designed for specific tasks, and additional effort is required for multi-task support. In contrast, LocoMan adopts a unified control framework, which uses the same low-level whole-body controller [43] for all five operation modes. By tracking desired trajectories at both the kinematics and dynamics level, our unified framework requires little task-specific tuning and performs all manipulation tasks at high accuracy.
@@ III Design of loco-manipulator
The design of LocoMan is inspired by the anatomy of human arms. As illustrated in Fig. 2(a), we attach custom-designed loco-manipulators to the calf of the front legs of a quadrupedal robot, and utilize the existing leg joints ($q_{1-3}$) together with the gripper joints ($q_{4-6}$) for 6DoF space manipulation. The original leg joints, $q_{1-3}$, similar to human shoulders and elbows, are primarily used for movement and positional tracking. The added joints of loco-manipulator, ($q_{4-6}$), similar to human wrist, are primarily used for orientation tracking. Integrated together, each front limb of LocoMan can reach a large variety of space poses effectively.
We carefully orient the joints of loco-manipulator to seamlessly integrate with the existing structure of the original robot. Specifically, we design the first joint ($q_{4}$) to be aligned with the calf joint of the front leg, design the last joint ($q_{6}$) to point towards the gripper, and design the second joint ($q_{5}$) to be perpendicular to the other two joints. In this way, the gripper can reach out to far spaces during manipulation (front-right foot of Fig. 2(a)), while stay tucked to the calf during standing and walking (front-left foot of Fig. 2(a)) with minimum interference to the locomotion capabilities.
@@ IV-B Overview
Effective loco-manipulation requires the robot to complete a large variety of tasks accurately and robustly, from in-place picking to locomotion while carrying objects. In light of this requirement, we design a unified control framework for LocoMan (Fig. 3), where a robot commander specifies the desired state of LocoMan based on the action command from user, and an unified whole-body controller (WBC) tracks the desired state based on LocoMan’s kinematics and dynamics model. The robot commander supports a variety of operation modes (Table III), such as locomotion and in-place manipulation, where the mode switch is governed by a finite state machine (FSM). Within each mode, the FSM launches a mode-specific mode commander to convert user specified target (e.g. walking speed or gripper pose) into desired states of the torso $(\overline{\bm{x}},\overline{\dot{\bm{x}}})^{\text{torso}}$, end effector $(\overline{\bm{x}},\overline{\dot{\bm{x}}})^{\text{eef}}$, and foot $(\overline{\bm{x}},\overline{\dot{\bm{x}}},\overline{\bm{c}})^{\text{foot}}$. We further design a set of mode switchers to smoothly transition between different operation modes. For robust operation and accurate tracking of the desired state, we extend a whole-body controller to solve the desired position $\overline{\bm{q}_{j}}$, velocity $\overline{\dot{\bm{q}}_{j}}$ and torque $\overline{\bm{\tau}_{j}}$ of each joint based on the complete kinematics and dynamics model of the robot. Then a joint-level PD controller computes the final torque $\bm{\tau_{j}}$ for each joint at a higher frequency. In addition, we implement two state estimators to accurately estimate robot state based on sensor feedback.
@@ IV-C Robot Commander
The robot commander supports five different operation modes (Table III), namely locomotion, locomanipulation, single-foot manipulation, single-gripper manipulation, and bimanual manipulation. Each mode takes in different user commands from either predefined planners, human motions, or joysticks, which are converted to desired states by a corresponding Task Commander.
For single-arm manipulation, the user specifies the desired torso 6D pose, and the desired gripper 6D pose or foot 3D position. For bimanual manipulation, the user commands the desired 6D poses for each gripper. The stance feet remain fixed to the ground during the three manipulation modes. For locomotion and loco-manipulation, the user specifies the desired velocity, roll, pitch, and height of the robot torso, together with the desired gripper orientation for the loco-manipulation mode.
To ensure smooth transition between operation modes, the FSM launches a specific mode switcher during mode transitions. For example, from locomotion mode to single-arm manipulation mode, LocoMan is commanded to zero velocity until all feet are on the ground, and then moves the torso to the rear opposite the end effector being manipulated.
@@ IV-D Unified Whole-Body Controller
Our WBC implementation is based on the two-step implementation from Kim et al. [43] with additional modeling of the loco-manipulators. In the first step, WBC tracks the desired states kinematically by computing the desired position $\overline{\bm{q}_{j}}$ and velocity $\overline{\dot{\bm{q}}_{j}}$ of each joint using Inverse Kinematics (IK). Since the robot may not have sufficient degrees of freedom to track the entire desired state, we assign a objective priority to each component of the desired states (Table III), and solve objectives in descending priority by projecting low-priority objectives into the null-space of high-priority ones. In the second step, WBC tracks the desired state dynamically by optimizing for joint torque commands $\overline{\bm{\tau}_{j}}$, where the objective is to track the desired torso acceleration, and the constraints include rigid body dynamics and friction cone limits. Please refer to the original work [43] for additional details.
@@ V Experiments
We design a set of experiments to validate the capability of LocoMan in versatile manipulation tasks. Specifically, we aim to answer the following questions: (1) How large is the workspace of LocoMan and how does it compare with the workspace of the original robot? (2) How well can LocoMan track a desired trajectories in its workspace? (3) What real-life manipulation tasks can LocoMan perform? and (4) What is the benefit of the leg-mounted manipulator design of LocoMan, compared with top-mounted arms [18, 21]?
@@ V-C Trajectory Tracking
Figure 5: The left figure shows trajectory tracking of the end effector position and the torso orientation when drawing a spatial square. The right figure depicts the corresponding 3D trajectory tracked by an LED light attached to the gripper.
@@ V-D Low-cost Vision-based Teleoperation Platform
In LocoMan, the robot torso and end effectors need to be commanded simultaneously, which would be difficult to teleoperate using a single joystick, keyboard or sensor glove. Therefore, we develop a low-cost vision-based teleoperation platform to teleoperate LocoMan using motions of human hands and torso. As illustrated in  Fig. 6 (a), the core of our teleoperation platform is a pair of 3D-printed rigid shells worn by palms and thumbs of both hands as wells as human torso. The shells are densely covered with AprilTags [44] whose corners and IDs can be robustly detected. Using a method similar to [45, 46], we set up calibrated high-resolution ZED cameras at multiple views and the 6D pose of the rigid shells can be obtained by minimizing the re-projection errors of AprilTag corners in the multiple-view images. The 6D pose of the torso and palm shells are directly used to command the 6D pose of the torso and end effector of quadrupedal robot, while the relative angles between the 6D poses of the thumb and palm shells are used to command the opening and closing of the robot gripper. In this way, the teleoperation users can move their torso and both of their hands and fingers to naturally control all motion of the quadrupedal robot. With a few cameras and mounts, our teleoperation platform is especially low-cost in hardware compared to previous works, while allowing the users to simultaneously command torso pose, two end effector poses and two gripper openings of the quadrupedal robot.
Figure 6: Low-cost teleoperation platform. (a) The rigid shell worn by palms, thumbs and torso created by 3D printing. (b) Simultaneous control of the 6D pose of the torso and the end effector of the quadrupedal robot. (c) Visualization of the estimated corresponding 6D poses of the user’s right palm, right thumb and the torso using 3D point clouds of AprilTags.

### v2

@@ Abstract
Quadrupedal robots have emerged as versatile agents capable of locomoting and manipulating in complex environments. Traditional designs typically rely on the robot’s inherent body parts or incorporate top-mounted arms for manipulation tasks. However, these configurations may limit the robot’s operational dexterity, efficiency, and adaptability, particularly in cluttered or constrained spaces. In this work, we present LocoMan, a dexterous quadrupedal robot with a novel morphology to perform versatile manipulation in diverse constrained environments. By equipping a Unitree Go1 robot with two low-cost and lightweight modular 3-DoF loco-manipulators on its front calves, LocoMan leverages the combined mobility and functionality of the legs and grippers for complex manipulation tasks that require precise 6D positioning of the end effector in a wide workspace. To harness the loco-manipulation capabilities of LocoMan, we introduce a unified control framework that extends the whole-body controller (WBC) to integrate the dynamics of loco-manipulators. Through experiments, we validate that the proposed whole-body controller can accurately and stably follow desired 6D trajectories of the end effector and torso, which, when combined with the large workspace from our design, facilitates a diverse set of challenging dexterous loco-manipulation tasks in confined spaces, such as opening doors, plugging into sockets, picking objects in narrow and low-lying spaces, and bimanual manipulation.
@@ I Introduction
Recent advances in the capability of quadrupedal robots have enabled them to traverse highly complex terrains [1, 2, 3, 4, 5, 6], perform acrobatic skills [7, 8, 9, 10, 11], and interact with humans [12]. While most of these works focus on improving the mobility of quadrupedal robots, the manipulation capability of quadrupedal robots still remains limited and often requires special designs to achieve tasks such as pressing buttons[13] and kicking balls[14, 15]. To be effectively deployed in daily life, quadrupedal robots are required to possess versatile manipulation skills with enhanced dexterity, in addition to their inherent locomotion capability.
Integrating manipulation into quadrupedal robots proves to be a difficult task. As illustrated in Table. I, without external hardware modifications, quadrupedal robots have to utilize their legs, head, or torso to move or transport objects [13, 15, 16, 17], which limits their ability to control the 6D poses of the object and perform high-precision tasks. A popular solution to address this challenge is to use a top-mounted robot arm [18]. However, this solution often comes at the cost of increased payload and decreased agility, and prevents the robot from reaching narrow spaces or performing more dexterous tasks such as bimanual manipulation. While foot-mounted grippers offer a lightweight alternative to full robot arms, existing work adopts 1-DoF grippers with limited functionality [19]. Therefore, it can be challenging to achieve versatile dexterous loco-manipulation using existing manipulator solutions for legged robots.
In this work, we present LocoMan, a comprehensive solution for versatile and skilled manipulation with legged robots. To reduce the payload and increase dexterity, LocoMan mounts two custom-designed lightweight 3-DoF manipulators on the front calves of the robot, and uses existing leg joints in addition to the manipulators for manipulation, as illustrated in Fig. 1. We refer to our designed manipulator as “loco-manipulator”, which draws conceptual inspiration from the front limbs of animals such as great apes and bears, which are adept at both locomotion and skilled manipulation.
@@
    We demonstrate that LocoMan can perform versatile real-life manipulation tasks, including cabinet opening, charger inserting, drink pouring, grasping in a narrow space, and basket hoisting.
@@ II-A Quadrupedal Loco-Manipulation
Without Additional Manipulators Researchers have a long history exploring the manipulation capability of quadrupedal robots. Without an additional manipulator, researchers have designed the robot to use its head [17, 22, 23, 24], torso [16], or foot [13, 14, 15, 25, 26, 27, 28] to interact with objects in the environment, where [22, 23, 24] rely on multiple robots to push larger objects. However, without an additional mechanism such as a gripper, these approaches are limited to relatively simple tasks that do not require fetching objects, such as kicking balls [14], pressing buttons [13], pushing boxes [24], opening doors [27], and probing [26].
With Additional Manipulators Many previous works mount a dedicated arm on top of the quadrupedal robot to perform more complicated manipulation tasks such as pick-and-place [29, 30], pulling doors [31, 21, 32], tablecloth spreading [33], turning wheels [34], wiping a whiteboard [18], and collaboratively carrying large objects [35, 36]. On the other hand, equipping a robot with a dedicated arm typically adds to the mechanical complexity and power requirements, leading to an increase in both weight and expense. Moreover, such an addition makes it difficult to operate in a narrow space and reduces the quadrupedal robot’s agility. Some other works alleviate these problems by installing more lightweight manipulators or grippers on the mouth [37] or foot [19, 20] of the quadrupedal robot. Tsvetkov et al. [19] propose a novel design of a small-scale quadrupedal robot with manipulators built into the legs, each requiring three additional actuators, where two are used for grasping. It facilitates single-arm and two-arm manipulation, but it only supports specifying target positions in joint space. Arm et al. [20] propose a learning-based controller that can track quadrupedal foot position targets through whole-body motions, and a 2-DoF gripper is attached to the foot to perform tasks such as collecting rock samples. However, how to efficiently reach a specified 6D pose of the end effector on the leg to perform more complex manipulation tasks is still underexplored. While some hexapod robots [38, 39, 40, 41, 42] use legs as manipulators, where integrated grippers can fetch objects with high degrees of freedom, prehensile manipulation with dexterity could be challenging for quadrupedal robots.
@@ II-B Controllers for Leg-Manipulation
To accomplish a wide range of manipulation tasks, the controller of the loco-manipulation robot need to support a variety of operation modes, including reaching a narrow space with a single arm, walking while carrying an object, and bi-manual manipulation with upright pose. While prior works have achieved similar tasks such as grasping [21, 31, 18], object pushing and kicking [17, 14, 15], and object carrying [41] using optimization-based [21, 31, 29, 17, 41] or learning-based [16, 14, 15, 18, 13] methods, most of these controllers are designed for specific tasks, and additional effort is required for multi-task support. In contrast, LocoMan adopts a unified control framework, which uses the same low-level whole-body controller [43] for all five operation modes. By tracking desired trajectories at both the kinematics and dynamics level, our unified framework requires little task-specific tuning and performs all manipulation tasks with high accuracy.
@@ III Design of loco-manipulator
The design of LocoMan is inspired by the anatomy of human arms. As illustrated in Fig. 2(a), we attach custom-designed loco-manipulators to the calf of the front legs of a quadrupedal robot, and utilize existing leg joints ($q_{1-3}$) together with the gripper joints ($q_{4-6}$) for 6DoF space manipulation. The original leg joints, $q_{1-3}$, similar to human shoulders and elbows, are used mainly for movement and positional tracking. The added joints of loco-manipulator, ($q_{4-6}$), similar to the human wrist, are primarily used for orientation tracking. Integrated together, each front limb of LocoMan can reach a wide variety of space poses effectively.
We carefully orient the joints of loco-manipulator to seamlessly integrate with the existing structure of the original robot. Specifically, we design the first joint ($q_{4}$) to be aligned with the calf joint of the front leg, design the last joint ($q_{6}$) to point towards the gripper, and design the second joint ($q_{5}$) to be perpendicular to the other two joints. In this way, the gripper can reach out to far spaces during manipulation (front-right foot of Fig. 2(a)), while staying tucked to the calf during standing and walking (front-left foot of Fig. 2(a)) with minimal interference with locomotion capabilities.
@@ IV-B Overview
Effective loco-manipulation requires the robot to complete a large variety of tasks accurately and robustly, from in-place picking to locomotion while carrying objects. In light of this requirement, we design a unified control framework for LocoMan (Fig. 3), where a robot commander specifies the desired state of LocoMan based on the action command from user, and a unified whole-body controller (WBC) tracks the desired state based on LocoMan’s kinematics and dynamics model. The robot commander supports a variety of operation modes (Table III), such as locomotion and in-place manipulation, where the mode switch is governed by a finite state machine (FSM). Within each mode, the FSM launches a mode-specific mode commander to convert user specified target (e.g. walking speed or gripper pose) into desired states of the torso $(\overline{\bm{x}},\overline{\dot{\bm{x}}})^{\text{torso}}$, end effector $(\overline{\bm{x}},\overline{\dot{\bm{x}}})^{\text{eef}}$, and foot $(\overline{\bm{x}},\overline{\dot{\bm{x}}},\overline{\bm{c}})^{\text{foot}}$. We further design a set of mode switchers to smoothly transition between different operation modes. For robust operation and accurate tracking of the desired state, we extend a whole-body controller to solve the desired position $\overline{\bm{q}_{j}}$, velocity $\overline{\dot{\bm{q}}_{j}}$ and torque $\overline{\bm{\tau}_{j}}$ of each joint based on the complete kinematics and dynamics model of the robot. Then a joint-level PD controller computes the final torque $\bm{\tau_{j}}$ for each joint at a higher frequency. In addition, we implement two state estimators to accurately estimate the robot state based on sensor feedback.
@@ IV-C Robot Commander
The robot commander supports five different operation modes (Table III), namely locomotion, locomanipulation, single-foot manipulation, single-gripper manipulation, and bimanual manipulation. Each mode commander converts different user commands from planners, human motions, or joysticks to desired states of LocoMan.
For single-arm manipulation, the user specifies the desired torso 6D pose, and the desired gripper 6D pose or foot 3D position. For bimanual manipulation, the user commands the desired 6D poses for each gripper. The stance feet remain fixed to the ground during the three manipulation modes. For locomotion and loco-manipulation, the user specifies the desired velocity, roll, pitch, and height of the robot torso, together with the desired gripper orientation for the later mode.
To ensure a smooth transition between operation modes, the FSM launches a specific mode switcher during mode transitions. For example, from locomotion mode to single-arm manipulation mode, LocoMan is commanded to zero velocity until all feet are on the ground, and then moves the torso to the rear opposite the end effector being manipulated.
@@ IV-D Unified Whole-Body Controller
Our WBC implementation is based on the two-step implementation from Kim et al. [43] with additional modeling of the loco-manipulators. In the first step, WBC tracks the desired states kinematically by computing the desired position $\overline{\bm{q}_{j}}$ and velocity $\overline{\dot{\bm{q}}_{j}}$ of each joint using Inverse Kinematics (IK). Since the robot may not have sufficient degrees of freedom to track the entire desired state, we assign a objective priority to each component of the desired states (Table III), and solve objectives in descending priority by projecting low-priority objectives into the null-space of high-priority ones. In the second step, WBC tracks the desired state dynamically by optimizing joint torque commands $\overline{\bm{\tau}_{j}}$, where the objective is to track the desired torso acceleration, and the constraints include rigid body dynamics and friction cone limits. Please refer to the original work [43] for additional details.
@@ V Experiments
We design a set of experiments to validate the capability of LocoMan in versatile manipulation tasks. Specifically, we aim to answer the following questions: (1) How large is the workspace of LocoMan and how does it compare with the workspace of the original robot? (2) How well can LocoMan track a desired trajectory in its workspace? (3) What real-life manipulation tasks can LocoMan perform? and (4) What is the benefit of the leg-mounted manipulator design of LocoMan, compared to the top-mounted arms [18, 21]?
@@ V-C Trajectory Tracking
Figure 5: The figure on the left shows trajectory tracking of the end effector position and the torso orientation when drawing a spatial square. The figure on the right depicts the corresponding 3D trajectory tracked by an LED light attached to the gripper.
@@ V-D Low-cost Vision-based Teleoperation Platform
In LocoMan, the robot torso and end effectors need to be commanded simultaneously, which would be difficult to teleoperate using a single joystick, keyboard or sensor glove. Therefore, we develop a low-cost vision-based teleoperation platform to teleoperate LocoMan using motions of human hands and torso. As illustrated in  Fig. 6 (a), the core of our teleoperation platform is a pair of 3D-printed rigid shells worn by palms and thumbs of both hands, as well as the human torso. The shells are densely covered with AprilTags [44] whose corners and IDs can be robustly detected. Using a method similar to [45, 46], we set up calibrated high-resolution ZED cameras at multiple views and the 6D pose of the rigid shells can be obtained by minimizing the re-projection errors of AprilTag corners in the multiple-view images. The 6D pose of the torso and palm shells are directly used to command the 6D pose of the torso and end effector of quadrupedal robot, while the relative angles between the 6D poses of the thumb and palm shells are used to command the opening and closing of the robot gripper. In this way, the teleoperation users can move their torso and both of their hands and fingers to naturally control all motion of the quadrupedal robot. With a few cameras and mounts, our teleoperation platform is especially low-cost in hardware compared to previous works, while allowing the users to simultaneously command torso pose, two end effector poses, and two gripper openings of the quadrupedal robot.
Figure 6: Low-cost teleoperation platform. (a) The rigid shells worn by palms, thumbs and torso created by 3D printing. (b) Simultaneous control of the 6D pose of the torso and the end effector of the quadrupedal robot. (c) Visualization of the estimated corresponding 6D poses of the user’s right palm, right thumb and the torso using 3D point clouds of AprilTag edges.